OpenStudy (anonymous):

divide the following radical and simplify the answer. 8/√9+√2 a. 8√3-8√2 b. 8/√3 c. 8 d. 8√3+8√2

OpenStudy (mertsj):

\[\frac{8}{\sqrt{9}+\sqrt{2}}\]

OpenStudy (mertsj):

\[\frac{8}{3+\sqrt{2}}\]

OpenStudy (mertsj):

Multiply numerator and denominator by 3-sqrt2

OpenStudy (anonymous):

its either a or d right?

OpenStudy (mertsj):